### v1

# LEKA: LLM-Enhanced Knowledge Augmentation

Abstract
Humans excel in analogical learning and knowledge transfer and, more importantly, possess a unique understanding of identifying appropriate sources of knowledge. From a model’s perspective, this presents an interesting challenge. If models could autonomously retrieve knowledge useful for transfer or decision-making to solve problems, they would transition from passively acquiring to actively accessing and learning from knowledge. However, filling models with knowledge is relatively straightforward—it simply requires more training and accessible knowledge bases. The more complex task is teaching models about which knowledge can be analogized and transferred. Therefore, we design a knowledge augmentation method LEKA for knowledge transfer that actively searches for suitable knowledge sources that can enrich the target domain’s knowledge. This LEKA method extracts key information from textual information from the target domain, retrieves pertinent data from external data libraries, and harmonizes retrieved data with the target domain data in feature space and marginal probability measures. We validate the effectiveness of our approach through extensive experiments across various domains and demonstrate significant improvements over traditional methods in reducing computational costs, automating data alignment, and optimizing transfer learning outcomes.
Introduction
Humans are good at identifying proper sources of knowledge. This is an ability that is rooted in our capability for analogical reasoning and knowledge management. On the contrary, artificial intelligence models do not inherently possess this intuition: they require explicit instructions and systematic training to identify and utilize relevant information. This gap presents a challenge in enriching domain knowledge and enhancing data augmentation.
Knowledge augmentation is vital for improving model performance, especially in domains with limited information or complex data structures (Tang et al. 2020). In knowledge augmentation, knowledge transfer is an important method that improves learning performance by transferring knowledge from external information sources. In data-limited scenarios, effective knowledge sourcing can bridge domain information gaps and enhance model robustness by leveraging relevant information from external sources. In these scenarios, knowledge augmentation, especially knowledge transfer can reduce reliance on extensive target domain data by strategically selecting sources matching the target needs.
Despite its potential, traditional knowledge augmentation methods often involve manual intervention, where human experts pick source domains based on their subjective interpretation of the target domain’s requirements. Using tabular learning as an example highlights several challenges: (1) structural and format differences across domains hinder data alignment and integration; (2) discrepancies in tasks and content between the selected source and target domains can diminish the success of knowledge augmentation; (3) extensive data preprocessing is often required to match the chosen source domain dataset properly. Additionally, relying on human expertise for source domain selection can lead to biases and inefficiencies, as these choices are typically made based on prior knowledge rather than through a rigorous, data-driven analysis.
A natural idea is to construct an automated search for datasets with relevant knowledge in a database with a similar structure to the target data, such as utilizing Retrieval-Augmented Generation (RAG). RAG combines retrieval systems with generative language models, and it enhances the model’s capabilities by providing access to an external library (Guo et al. 2017; Song et al. 2019; Shi et al. 2018). For dataset retrieval, some existing works (Fleischer et al. 2024; Siriwardhana et al. 2023) use the retrieved documents as contextual information for generation and significantly enrich the model’s knowledge base.
However, this direct method often proves impractical due to the high computational costs of embedding entire domains or whole datasets. Additionally, the retrieved knowledge may not align effectively with the intended application. Such misalignments typically result from variations in data distribution, feature spaces, or contextual differences between the source and target data. Furthermore, the quality of the retrieved data, such as the noise or incomplete features, can also hinder its effectiveness for augmenting data via knowledge augmentation. Thus, in such situations, identifying the optimal knowledge source that provides the most relevant and high-quality data tailored to the specific needs of the target data through the capabilities of LLMs remains a substantial challenge.
Figure 1: Example of LEKA. We adopt an LLM to retrieve proper source domain data to transfer knowledge to a data-limited target domain. The LLM extracts the key information of target data to retrieve for a relevant dataset; then we adopt the LLM for harmonization.
Our Targets. Comprehensively, we aim to address three main challenges in retrieving source domain data: (1) how to effectively extract key information based on the target domain with limited computational costs; (2) how to design an efficient automatic source domain retrieval for knowledge augmentation; and (3), how to automatically harmonize the retrieved source data with the target domain to optimize transfer efficiency and improve learning outcomes.
Our Approach. To address the challenges above, we design LLM-Enhanced Knowledge Augmentation (LEKA), a novel and automated data retrieval and augmentation method by knowledge augmentation. Specifically, (1) we utilize an LLM to extract the key textual information of the target domain; (2) we deploy dataset RAG in an external database to efficiently extract relevant knowledge to target domain dataset. The RAG libraries can contain large amounts of continuously updated datasets, ensuring that the data retrieved from these libraries is more precise and timely; (3) then the LLM automatically harmonizes the retrieved source data with the target domain on feature space and marginal probability measures to enhance downstream task learning performance. The data harmonization reduces the structural and semantic differences between the source and target data. This data harmonization is crucial for reducing domain shift, which optimizes transfer efficiency and improves the overall learning outcomes of the model.
As shown in Figure 1, for a target domain targeting predicting a rare cancer, direct learning is challenging due to insufficient sample information. To augment the data with external information, we need human experts with knowledge of both cancer physiological indicators and machine learning to retrieve datasets. However, our method can automate the retrieval of a common cancer dataset as the knowledge source with extensive feature information. This retrieved dataset has a more complete and transparent feature space with a structure similar to the target domain. The LLM then adjusts the retrieved data based on the characteristics of the target domain. In this way, we achieve a highly harmonized source dataset for knowledge augmentation.
In summary, our contribution includes:
  1. 1.
    We introduce a novel automated data augmentation method, LEKA, which utilizes an automated retrieval method for external data and harmonizes it with target data for knowledge augmentation.
  2. 2.
    We develop a novel paradigm by incorporating an LLM into the data harmonization process to optimize data space and structure to enhance downstream machine learning performance.
  3. 3.
    We conduct a series of experiments to validate the effectiveness and robustness of our LEKA method across different tasks. Experimental results demonstrate that our method has clear advantages over existing methods.
Related Work
Knowledge Transfer
Knowledge transfer is a knowledge augmentation method that improves learning on new tasks by transferring knowledge from a related task (Alyafeai, AlShaibani, and Ahmad 2020; Iman, Arabnia, and Rasheed 2023; Bu et al. 2021). Knowledge transfer allows cross-domain knowledge transformation despite data distribution or feature space differences. Specifically, knowledge transfer adapts models developed for one task to perform better on different but related tasks, as it adjusts the feature mappings and decision boundaries to suit new knowledge (He et al. 2021; Han, Pang, and Wu 2021; Yordanov et al. 2021). Finding suitable pre-training knowledge is crucial for knowledge transfer because it significantly enhances the effectiveness of the model by providing a relevant starting point. While many methods exist to adjust existing knowledge within the same or across different domains to improve transfer learning outcomes, the process of retrieving certain knowledge still heavily relies on manual effort.
Data Harmonizing with LLMs
Data harmonization using LLMs offers significant benefits by leveraging their natural language capabilities to standardize and integrate diverse datasets and further enhance model performance through improved data consistency (Cao, Gorodetsky, and Mitkas 2009; Liu et al. 2021; Durante et al. 2024; Zhang et al. 2024d). However, this approach has several challenges, including the high computational costs of training LLMs and potential biases in training data, which can adversely affect the harmonization process (Feng et al. 2021; Li, Hou, and Che 2022; Bu et al. 2019). Furthermore, the risk of overfitting remains a concern, as models may become overly specialized in training data nuances, reducing their effectiveness on new datasets.
Retrieval Augmented Generation for Knowledge Augmentation
Retrieval Augmented Generation (RAG) (Lewis et al. 2020) for knowledge augmentation is a method that enhances the capabilities of generative language models by integrating information retrieval with model generation (Hu and Lu 2024). This technique aids LLMs in tasks that demand deep and specific domain knowledge (Zhang et al. 2024b; Huang and Huang 2024; Zhang et al. 2024a). RAG for knowledge augmentation provides access to expansive external libraries (collections of documents or domain-relevant knowledge), making it especially effective for transferring knowledge across different domains (Siriwardhana et al. 2023; Dan, Lu, and Li 2024). RAG for knowledge augmentation embeds extensive databases directly into the generative process so that specific, domain-related information is both accessible and effectively utilized. In this way it can significantly enhance the performance for complex tasks in knowledge augmentation scenarios.
Preliminary
Definition
Definition 1. (Domain)
A domain $\mathcal{D}$ is an ordered pair consisting of a feature space $\mathcal{X}$ and a marginal probability measure $P$ defined on this feature space. In other words, $\mathcal{D}=(X,P)$, where $\mathbf{X}=\{\mathbf{x}|\mathbf{x}_{i}\in X,i=1,\ldots,n\}$ is an instance set. And $P$ is a probability measure that describes the probability of occurrence of feature vectors $\mathbf{x}\in\mathbf{X}$. This probability measure $P$ makes $(\mathcal{X},\mathcal{B}(\mathcal{X}),P)$ a probability space, where $\mathcal{B}(\mathcal{X})$ is the Borel $\sigma$-algebra on $\mathcal{X}$.
Source Domain. The source domain $\mathcal{D}_{S}$ consists of instances paired with labels $y$, represented as $\mathcal{D}_{S}=\{(\mathbf{x},y)|\mathbf{x}_{i}\in\mathcal{X}_{S},y_{i}\in\mathcal{Y}_{S},i=1,\ldots,n^{S}\}$. Here, $\mathcal{X}_{S}$ is the feature space and $\mathcal{Y}_{S}$ is the label space for the source domain. This domain provides labeled data used to train models in preparation for transfer learning tasks.
Target Domain. The target domain $\mathcal{D}_{T}$ typically contains a mix of unlabeled instances and a smaller set of labeled instances, denoted as $\mathcal{D}_{T}=\{\mathbf{x}\in\mathcal{X}_{T}\}\cup\{(\mathbf{x},y)|(\mathbf{x},y)\in\mathcal{X}_{T}\times\mathcal{Y}_{T}\}$. Here, $\mathcal{X}_{T}$ is the feature space and $\mathcal{Y}_{T}$ is the label space for the target domain. We aim to evaluate and fine-tune the transfer learning models with data in the target domain.
Definition 2. (Task)
A task $\mathcal{T}$ consists of a label space $\mathcal{Y}$ and a decision function $f$, formally noted as $\mathcal{T}=(\mathcal{Y},f)$, where $\mathcal{Y}$ is a metric space that contains all possible labels, and $f$ is a mapping from the feature space $\mathcal{X}$ to a set of conditional probability measures on the label space $\mathcal{Y}$.
Source Task. The learning task of the source task $\mathcal{T}_{S}$ is typically represented as learning a target function $f_{S}:\mathcal{X}_{S}\to\mathcal{Y}_{S}$, where $\mathcal{Y}_{S}$ is the label space of the source task.
Target Task. The learning task of the target task $\mathcal{T}_{T}$ is typically represented as learning a target function $f_{T}:\mathcal{X}_{T}\to\mathcal{Y}_{T}$, where $\mathcal{Y}_{T}$ is the label space of the target task.
Definition 3. (Knowledge Transfer)
Knowledge transfer is an augmentation method that adopts observations from source domains and tasks, denoted as $\{(\mathcal{D}_{S_{i}},\mathcal{T}_{S_{i}})|i=1,\ldots,m^{S}\}$. Here $m^{S}\in\mathbb{N}^{+}$ represents the number of source domains and tasks. Similar observations from target domains and tasks are denoted as $\{(\mathcal{D}_{T_{j}},\mathcal{T}_{T_{j}})|j=1,\ldots,m^{T}\}$, where $m^{T}\in\mathbb{N}^{+}$. The goal of knowledge transfer is to utilize the knowledge embedded in the source domains $\mathcal{D}_{S_{i}}$ to enhance the performance of the learned decision functions $f_{T_{j}}$ across the target domains $\mathcal{D}_{T_{j}}$, for $j=1,\ldots,m^{T}$.
Problem Formulation
We formulate the task to enhance the performance of knowledge transfer through automated data retrieval and harmonization using an LLM. Concretely, we improve the performance of the learned decision function $f_{T_{j}}$ by reconstructing and refining the source domain $\mathcal{D}_{S}$. In this way, we better utilize the knowledge implied in $\mathcal{D}_{S}$ as we mitigate domain shifts and facilitate a more effective transfer of learned models. Concretely, our optimization objective is to retrieve and reconstruct a source domain $\mathcal{D}_{S}^{*}$:
$\mathcal{D}_{S}^{*}=\underset{\hat{\mathcal{D}_{S}}}{\text{argmax}}\;\mathbf{\mathcal{P}}_{\mathcal{D}_{T}}(f_{T_{j}}),$ (1)
where $\mathcal{P}$ is the performance indicator of $f_{T_{j}}$ and $\hat{\mathcal{D}_{S}}$ is a reconstructed source domain aligned with $\mathcal{D}_{T}$.
Methodology
Figure 2: Framework of LEKA includes: 1) an LLM extracts and embeds the textual information of target dataset, then 2) retrieves datasets in libraries and 3) process data harmonization. With harmonized dataset, we can transfer knowledge from the source dataset we construct to enhance target dataset learning.
In this section, we introduce our novel knowledge augmentation method named LLM-Enhanced Knowledge Augmentation (LEKA), designed to dynamically and automatically retrieve and refine source data to transfer knowledge and further enhance target data learning. Specifically, our study focuses on tabular datasets as the concrete form of data. The LEKA leverages an LLM to extract essential information from the target dataset, including data structure, feature names, and descriptions. Within this framework, the LLM uses its capability to analyze and synthesize data to optimize the selection and refinement of source datasets. It extracts and summarizes keywords for retrieval tailored to the target domain and task. After retrieval, the LLM refines and harmonizes these datasets in feature space and marginal probability measures with the target dataset. Figure 2 illustrates the overview of the LEKA framework, which comprises three stages: (1) Dataset Retrieval; (2) Data Harmonization; and (3) Knowledge Transfer and Evaluation.
Dataset Retrieval
In this phase, our primary objective is to identify and retrieve a source dataset $\mathcal{D}_{S}$ similar to the target dataset $\mathcal{D}_{T}$ for effective pre-training in knowledge transfer scenarios. To achieve this, we utilize an LLM $\mathcal{A}$ to analyze and retrieve data that matches the structural properties and purposes of $\mathcal{D}_{T}$. A straightforward method is to embed the target dataset $\mathcal{D}_{T}$ by $\mathcal{A}$ to capture its essential features. However, this approach can be computationally intensive and prone to errors. Inspired by the strategies outlined in (Zhang et al. 2024c), we turn to leverage the textual information, including the data descriptions $\mathcal{C}_{T}$ and feature names $l_{T}$ of $\mathcal{D}_{T}$. These textual elements provide rich insights into the structure and purpose of $\mathcal{D}_{T}$, and the LLM can precisely focuses on the semantic content of $\mathcal{D}_{T}$.
First, the LLM $\mathcal{A}$ constructs a query $\mathcal{Q}$ based on this textual information:
$\mathcal{Q}=\mathcal{A}(\mathcal{C}_{T},l_{T}),$ (2)
and then embeds the query:
$q=\mathcal{A}(\text{embed}(\mathcal{Q})),$ (3)
where $\text{embed}(\cdot)$ is the embedding process that transforms the input into a vector representation.
We then retrieve through in a library $\mathcal{L}$ of dataset for the top-$k$ most relevant datasets $\{\mathcal{D}_{k}\}$ to query $\mathcal{Q}_{t}$. We evaluate the relevance of all datasets in $\mathcal{L}$ to the query $\mathcal{Q}$ by the cosine similarity of the embeddings and select the top-$k$ datasets with the highest similarity:
$\text{sim}(q,d_{k})=\frac{q\cdot d_{k}}{\|q\|\|d_{k}\|},$ (4)
where $d_{k}=\mathcal{A}(\text{embed}(\mathcal{D}_{k}))$ is the embedding of a potential dataset $\mathcal{D}_{k}$’s textual information, and $\|\cdot\|\in\mathbb{R}$ is the norm function.
In this phase, we map textual information to a high-dimensional vector space. Essentially, we approximate the probability distributions of the datasets in the library to retrieve $k$ source datasets most similar in structure and purpose to the target dataset. We leverage the geometric properties of vectors in high-dimensional space to assess and quantify dataset similarities. This prepares us for further dataset processing for transfer learning.
Data Harmonization
Now we have $k$ potential source domain datasets, we turn to align these datasets with the target dataset $\mathcal{D}_{T}$ with an LLM. This alignment process is to adjust the feature space and marginal probability measures of the source dataset $\mathcal{D}_{k}$ to closely match those of the target dataset $\mathcal{D}_{T}$. This process involves transformations of features and adjustments of distributions. Here we take $\mathcal{D}_{k}$ as an example.
Feature Space Transformation.
We denote $\mathcal{X}_{k}$ and $\mathcal{X}_{T}$ to be the feature spaces of the source dataset $\mathcal{D}_{k}$ and the target dataset $\mathcal{D}_{T}$, respectively. Our LLM constructs a mapping function $f_{map}:\mathcal{X}_{k}\rightarrow\mathcal{X}_{T}$ for alignment. This function transforms the features in $\mathcal{X}_{k}$ to minimize the distance between the transformed source features and the target features. Here, we aim to minimize the distance $d(f(\mathcal{X}_{k}),\mathcal{X}_{T})$, where $d$ is a kernel distance:
$d(f(\mathcal{X}_{k}),\mathcal{X}_{T})=\sqrt{\sum_{i,j}(\kappa(f(x_{i}^{k}),f(x_{j}^{k}))-\kappa(x_{i}^{T},x_{j}^{T}))^{2}}.$ (5)
Here, $\kappa(\cdot,\cdot)$ is a kernel function, $x_{i}^{k}$ and $x_{j}^{k}$ are the $i$-th and $j$-th feature vectors in the source dataset $\mathcal{D}_{k}$. We adopt a Gaussian kernel $\kappa(x,y)=\exp(-\gamma\|x-y\|^{2})$, with $\gamma$ being a positive bandwidth parameter. This kernel function can effectively measure the similarity between pairs of points in high-dimensional spaces and captures both linear and non-linear relationships. Thus, it is capable of handling the complexities inherent in high-dimensional data. Specifically, compared with Euclidean distances, kernel distances can capture the geometric structure of the data manifold. It provides a more robust and informative measure of similarity for the source-target alignment.
Marginal Probability Measures Harmonization.
Our LLM then aligns the marginal probability measures $\mathcal{P}_{k}$ and $\mathcal{P}_{T}$ of the source and target datasets. The LLM analyzes and refines the textual information including labels, feature names, and classification probability distribution,.etc. To measure the differences between $\mathcal{D}_{k}$ and $\mathcal{D}_{T}$, we adopt the Wasserstein distance:
[…]
$\mathcal{D}_{k}^{{}^{\prime}}=\mathcal{A}(\mathcal{D}_{k},\mathcal{P}_{k},\mathcal{P}_{T}),$ (7)
where $\Gamma(\mathcal{P}_{k},\mathcal{P}_{T})$ represents the set of all joint distributions with marginals $\mathcal{P}_{k}$ and $\mathcal{P}_{T}$ on $\mathcal{X}\times\mathcal{X}$. The Wasserstein distance minimizes the transportation cost while preserving the geometric structure of the data distributions. Specifically, the Wasserstein distance calculates the minimum “geographical” cost required to move data from one distribution to another, where “geographical” cost refers to the cost of moving data from one point to another in the feature space.
Moreover, the Wasserstein distance is particularly useful when dealing with distributions whose support sets (i.e., the effective range or set of the distributions) do not fully overlap. In real-world data applications, it is common for the source and target datasets to originate from different distributions. Their data points do not align perfectly and cover the same areas. In these cases, traditional distance metrics like the Euclidean distance may not effectively reflect the true differences between the two datasets, as they merely measure differences in position but ignore the overall structure of the data distributions.
With this source-target alignment process, the LLM improves the efficiency of knowledge transfer from the source domain to the target domain. Thus it enhances the overall performance of our transfer learning models. Further, we optimize data handling and boost the adaptability and accuracy of transfer learning to be adopted in various applications.
Knowledge Transfer and Evaluation
After aligning the source dataset $\mathcal{D}_{k}^{\prime}$ with the target dataset $\mathcal{D}_{T}$, and adjusting the marginal probability measures $\mathcal{P}_{k}^{\prime}$ to $\mathcal{P}_{T}$, we proceed to integrate these aligned datasets into the transfer learning model. At this stage, our goal is to effectively transfer knowledge from $\mathcal{D}_{k}^{\prime}$ to $\mathcal{D}_{T}$ by minimizing a defined loss function while adapting the model to the target domain.
To achieve this, the transfer learning process focuses on updating the decision function $f_{T_{j}}$. Specifically, we aim to minimize the expected loss over the target dataset while incorporating the knowledge transferred from the source dataset:
$f_{T_{j}}^{*}=\arg\min_{f_{T_{j}}}\mathbb{E}_{(x,y)\in\mathcal{D}_{T}}[\mathcal{L}(f_{T_{j}}(x;\theta),y)],$ (8)
where $\mathcal{L}$ is the loss function, $x$ represents the features, $y$ represents the labels in $\mathcal{D}_{T}$, and $\theta$ denotes the parameters of $f_{T_{j}}$ that are being optimized. We calculate the expectation by the probability distribution $\mathcal{P}_{T}$, which has been closely aligned with $\mathcal{P}_{k}^{\prime}$ to ensure consistency and maximize the efficacy of the knowledge transfer.
During update, the optimization of model parameters $\theta$ leverages both the target dataset $\mathcal{D}_{T}$ and the aligned source dataset $\mathcal{D}_{k}^{\prime}$, which incorporates domain-specific characteristics to reduce domain shift. We define the optimization process as:
$\theta^{*}$ $=\arg\min_{\theta}\bigg{(}\alpha\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}[\mathcal{L}(f_{T_{j}}(x;\theta),y)]$ (9)
$\quad+(1-\alpha)\mathbb{E}_{(x,y)\in\mathcal{D}_{T}}[\mathcal{L}(f_{T_{j}}(x;\theta),y)]\bigg{)}.$ (10)
This formula aims to fine-tune the decision function $f_{T_{j}}$ by minimizing the weighted sum of expected losses across the datasets. The loss function $\mathcal{L}(f_{T_{j}}(x;\theta),y)$ evaluates prediction accuracy, guiding the adjustment of parameters. The expectations $\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}$ and $\mathbb{E}_{(x,y)\in\mathcal{D}_{T}}$ represent the mean losses over the source and target datasets, respectively. The weighting factor $\alpha$ adjusts the relative influence of each dataset, and keeps flexible adaptation between leveraging established knowledge from the source and integrating new data from the target. In this way, we utilize the transferred knowledge to enhance model performance on the target task.
Following the optimization of the model parameters, the backpropagation process then updates $\theta$ by minimizing the total loss. This loss is a weighted sum calculated from the losses on $\mathcal{D}_{k}^{\prime}$ and $\mathcal{D}_{T}$. This involves calculating the gradient of the loss function with respect to $\theta$ and updating $\theta$ using gradient descent methods:
$\theta\leftarrow\theta-\eta\nabla_{\theta}\left(\alpha\mathcal{L}(\mathcal{D}_{k}^{\prime};\theta)+(1-\alpha)\mathcal{L}(\mathcal{D}_{T};\theta)\right),$ (11)
where $\eta$ is the learning rate. The gradients are computed based on both datasets, which allows the model to learn both from the aligned source data and the target data.
After the parameter optimization and backpropagation processes, we evaluate the effectiveness of the transfer learning process; we adopt a performance metric $\phi$ on the target dataset:
$\phi(f_{T_{j}}^{*},\mathcal{D}_{T}),$ (12)
where $\phi$ measures the performance of the optimized model $f_{T_{j}}^{*}$ on $\mathcal{D}_{T}$.
In this way, we systematically improve the model based on empirical performance metrics.
Table 1: Datasets description. Here we use four datasets from medical and economic domains.
Datasets Samples Features Class
BCW 570 30 2
VID 8631 22 50
HD 303 13 2
TCC 7043 21 2
Experiments
In this section, we present four experiments to demonstrate the effectiveness and impacts of the LEKA. First, we compare the performance of the LEKA against several baseline methods on four downstream tasks. Second, we present the correlations between several target domains and their retrieved source domain. Finally, we discuss the reason for the improvement of performance.
Table 2: Performance comparison of transfer learning methods.
Dataset Metrics FTT TTab LEKA
BCW Acc 0.956 0.956 0.991
Prec 0.951 0.948 0.988
Rec 0.960 0.960 0.993
F1 0.955 0.954 0.990
VID Acc 0.745 0.797 0.995
Prec 0.758 0.588 0.996
Rec 0.747 0.513 0.996
F1 0.739 0.519 0.996
HD Acc 0.738 0.803 0.918
Prec 0.726 0.802 0.914
Rec 0.718 0.802 0.918
F1 0.721 0.802 0.916
TCC Acc 0.836 0.795 0.887
Prec 0.803 0.738 0.846
Rec 0.865 0.712 0.901
F1 0.817 0.722 0.865
Figure 3: Comparison of accuracy and F1 scores on various transfer learning methods.
Experiment Settings
Datasets and Domains. We evaluate our method primarily on four datasets of medical and economic domains: (1) Breast Cancer Wisconsin (Diagnostic) (BCW) (Wolberg et al. 1995), (2) Heart Disease (HD) (Janosi et al. 1988), (3) Vehicle Insurance Data (VID) (Bhatt 2019), and (4) Telco Customer Churn (TCC) (Zhang, Moro, and Ramos 2022). We show the detailed information about features of the datasets in Table 1.
Metrics and Models. We evaluate the model performance by the following metrics: Overall Accuracy (Acc) measures the proportion of true results (both true positives and true negatives) in the total dataset. Precision (Prec) reflects the ratio of true positive predictions to all positive predictions for each class. Recall (Rec), also known as sensitivity, reflects the ratio of true positive predictions to all actual positives for each class. F-Measure (F1) is the harmonic mean of precision and recall, providing a single score that balances both metrics. We apply the LEKA across a range of models: 1) Tabnet (TN) (Arik and Pfister 2021); 2) TabTransformer (TT) (Huang et al. 2020); 3) Random Forest (RF) (Rigatti 2017); 4) Gradient Boosting Decision Trees (GBDT) (Lin et al. 2023) ; 5) XGBoost (XB) (Chen and Guestrin 2016). We compare the performance outcomes in these tasks both with and without our method.
Baseline Models. We compare the LEKA method with four baseline methods, including: 1) Raw Data: direct using vanilla data; 2) TIFG (Zhang et al. 2024c): feature generation with LLM; 3) KPDDS (Huang et al. 2024): data synthesis with LLM using key points examples; 4) GReaT (Borisov et al. 2022): data synthesis with LLM simulating data subset distributions; 5) FTT (Levin et al. 2022) A TabTransformer pretrained on the source domain and then fine-tuned on the target domain; 6) TTab (Wang and Sun 2022) A transferable tabular Transformer capable of learning from multiple tabular datasets.
Implementation Details. In our setup for data synthesis and model training, we utilize GPT-4o as the query generator, combined with the Exa API to fetch web pages containing datasets from Kaggle and the UCI Machine Learning Repository that may be suitable for knowledge transfer. We extract dataset descriptions from web pages, and GPT-4o assesses their potential for knowledge transfer. Additionally, GPT-4o serves as a generator for executable code, performing up to five code generations. For our models, we configure TN, TT, and FTT with a batch size of $512$ for the VID and TCC datasets and a batch size of $32$ for the BCW dataset, a maximum of $100$ epochs, and employ early stopping with a patience of $20$. The learning rate is set at the default $0.02$ for pytorch_tabnet. For the RF and GBDT models, the number of trees is set to $100$, with GBDT also configured with a learning rate of $0.1$ and a max depth of $3$. TTab is set with a maximum epoch of $50$, a learning rate of $1\times 10^{-3}$, and a weight decay of $1\times 10^{-4}$.
Table 3: Overall performance on downstream tasks. The best results are highlighted in bold, and the runner-up results are highlighted in underline. (Higher values indicate better performance.)
Metrics Model GBDT RF TN XB TT
Acc BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC
Raw 0.939 0.991 0.754 0.793 0.921 0.827 0.770 0.776 0.956 0.436 0.803 0.783 0.930 0.946 0.770 0.773 0.965 0.714 0.574 0.754
KPDDS 0.947 0.992 0.787 0.787 0.939 0.846 0.803 0.787 0.982 0.567 0.820 0.806 0.956 0.948 0.836 0.789 0.974 0.728 0.639 0.781
GReaT 0.974 0.994 0.803 0.808 0.947 0.847 0.852 0.788 0.965 0.506 0.885 0.795 0.974 0.953 0.820 0.784 0.974 0.724 0.656 0.754
TIFG 0.947 0.994 0.836 0.815 0.947 0.858 0.869 0.800 0.982 0.579 0.902 0.815 0.965 0.952 0.852 0.803 0.965 0.736 0.705 0.790
LEKA 0.991 0.995 0.869 0.825 0.982 0.868 0.885 0.811 0.982 0.596 0.918 0.838 0.991 0.960 0.885 0.823 0.991 0.787 0.754 0.887
Prec BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC
Raw 0.948 0.992 0.804 0.734 0.926 0.834 0.751 0.716 0.949 0.429 0.803 0.717 0.928 0.932 0.775 0.706 0.965 0.742 0.302 0.709
KPDDS 0.936 0.993 0.782 0.720 0.929 0.850 0.807 0.725 0.978 0.594 0.823 0.767 0.954 0.934 0.843 0.744 0.974 0.752 0.817 0.724
GReaT 0.974 0.994 0.804 0.752 0.949 0.851 0.852 0.743 0.959 0.448 0.895 0.746 0.966 0.955 0.823 0.748 0.979 0.732 0.659 0.697
TIFG 0.947 0.994 0.840 0.767 0.940 0.863 0.867 0.750 0.981 0.571 0.900 0.748 0.967 0.957 0.849 0.750 0.965 0.725 0.707 0.734
LEKA 0.988 0.996 0.863 0.782 0.980 0.869 0.887 0.760 0.977 0.619 0.914 0.823 0.989 0.962 0.885 0.785 0.991 0.802 0.752 0.846
Rec BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC
Raw 0.912 0.991 0.770 0.701 0.918 0.828 0.751 0.671 0.954 0.443 0.803 0.680 0.928 0.926 0.769 0.682 0.965 0.720 0.461 0.757
KPDDS 0.959 0.993 0.785 0.689 0.951 0.845 0.806 0.675 0.986 0.553 0.819 0.686 0.962 0.930 0.845 0.710 0.974 0.734 0.522 0.755
GReaT 0.968 0.994 0.804 0.717 0.944 0.846 0.853 0.702 0.967 0.502 0.887 0.692 0.973 0.954 0.827 0.704 0.968 0.726 0.661 0.717
TIFG 0.935 0.994 0.839 0.725 0.940 0.856 0.867 0.697 0.981 0.570 0.904 0.700 0.957 0.951 0.852 0.711 0.965 0.733 0.690 0.724
LEKA 0.993 0.996 0.877 0.740 0.985 0.865 0.887 0.725 0.986 0.603 0.918 0.732 0.993 0.958 0.885 0.747 0.991 0.789 0.768 0.901
F1 BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC
Raw 0.927 0.991 0.750 0.714 0.920 0.827 0.751 0.685 0.952 0.400 0.803 0.693 0.928 0.927 0.769 0.692 0.965 0.704 0.365 0.718
KPDDS 0.944 0.993 0.783 0.701 0.936 0.844 0.803 0.691 0.982 0.510 0.819 0.708 0.956 0.930 0.836 0.723 0.974 0.724 0.429 0.735
GReaT 0.971 0.994 0.803 0.731 0.946 0.846 0.852 0.716 0.963 0.440 0.885 0.709 0.970 0.952 0.819 0.718 0.973 0.712 0.655 0.704
TIFG 0.941 0.994 0.836 0.741 0.940 0.853 0.867 0.715 0.981 0.535 0.901 0.718 0.961 0.951 0.850 0.726 0.965 0.706 0.691 0.729
LEKA 0.990 0.996 0.866 0.756 0.982 0.864 0.885 0.739 0.981 0.568 0.916 0.760 0.991 0.959 0.885 0.761 0.991 0.780 0.750 0.865
Experiment Results
Overall Performance. The experiment results are shown in Table 3, where we compare our LEKA method in four datasets. In summary:
(1) Compared with baseline models, LEKA significantly improves accuracy, surpassing baseline methods like TIFG and GreAT, with enhancements up to $4.4\%$ in models like GBDT. LEKA consistently achieves top results in precision and recall, indicating its precision in correctly identifying relevant cases while minimizing false positives. The F1 scores under LEKA are notably high, reflecting its effective balance between precision and recall across various models and datasets. These results validate the effectiveness of LEKA ’s retrieval and harmonization strategies and the robustness in diverse application scenarios. Overall, LEKA’s strategic approach to knowledge transfer is particularly advantageous in complex data environments, showcasing its adaptability and efficiency compared to traditional methods.
(2) The LEKA outperforms all baseline methods in almost all metrics and datasets. Specifically, LEKA shows enhancements in accuracy by $2-5\%$ over other methods. It demonstrates notable improvements of up to $4.4\%$ in GBDT for the VID dataset and $5.4\%$ in RF for the BCW dataset. For improvement of overall accuracy, we demonstrate the LEKA’s effective retrieval and harmonization of feature space and marginal probability measures with the target datasets. Meanwhile, the LEKA consistently outperforms baselines in reducing misclassification rates with the highest F1 scores. These results prove that LEKA’s retrieval and data harmonization reduce the likelihood of misclassification by forming a deeper and clearer understanding of the potential relationship between features.
Comparison with transfer learning methods. We then compare our LEKA method with transfer learning methods to demonstrate its effectiveness in enhancing model performance in complex domains. In these scenarios, the transfer learning methods show their reliance on manual data selection and alignment processes. The results demonstrate that LEKA outperforms transfer learning methods FFT and TTab across various metrics and datasets. For example, in the BCW dataset, LEKA improves accuracy by $3.5\%$ and recall by $3.3\%$ compared to the competing methods. This superior performance is consistent across other datasets like VID, HD, and TCC, with notable enhancements in precision and recall, emphasizing LEKA’s effective data harmonization capabilities. This adaptability and efficiency in handling diverse datasets underscore LEKA’s robustness and its significant advantages.
Conclusion
In this paper, we introduce LLM-Enhanced Knowledge Augmentation (LEKA), a novel retrieval and harmonization framework that dynamically refines source data for effective knowledge transfer. This structure significantly enhances the process of data augmentation by leveraging advanced LLM capabilities to automatically align and optimize data retrieved from diverse external libraries. Our extensive experiments across various tasks demonstrate the effectiveness and superiority over existing methods, especially in improving model adaptability and accuracy in complex data environments. By adopting LEKA on data-scarce domains, we achieve substantial improvements in learning performance and domain-specific task accuracy. For future work, we plan to extend this framework to include more varied data types and conduct a thorough empirical analysis to gain further insights into its underlying mechanisms and impacts.
Limitations
We acknowledge the following limitations of our work: (1) The current work is only tested on tabular data. More complex test scenarios are not involved; (2) Despite its automation, LEKA can be computationally intensive. This could limit its applicability in resource-constrained environments or require substantial computational resources to maintain operational efficiency; (3) Adopting the LEKA in tasks with unique requirements may have inherent limitations.

### v2

@@
 The more complex task is teaching models about which knowledge can be analogized and transferred.
-Therefore, we design a knowledge augmentation method LEKA for knowledge transfer that actively searches for suitable knowledge sources that can enrich the target domain’s knowledge.
-This LEKA method extracts key information from textual information from the target domain, retrieves pertinent data from external data libraries, and harmonizes retrieved data with the target domain data in feature space and marginal probability measures.
+Therefore, we design a knowledge augmentation method, LEKA, for knowledge transfer that actively searches for suitable knowledge sources that can enrich the target domain’s knowledge.
+This LEKA method extracts key information from the target domain’s textual information, retrieves pertinent data from external data libraries, and harmonizes retrieved data with the target domain data in feature space and marginal probability measures.
 We validate the effectiveness of our approach through extensive experiments across various domains and demonstrate significant improvements over traditional methods in reducing computational costs, automating data alignment, and optimizing transfer learning outcomes.
-Introduction
+1 Introduction
 Humans are good at identifying proper sources of knowledge.
 This is an ability that is rooted in our capability for analogical reasoning and knowledge management.
-On the contrary, artificial intelligence models do not inherently possess this intuition: they require explicit instructions and systematic training to identify and utilize relevant information.
+In contrast, artificial intelligence models do not inherently possess this intuition: they require explicit instructions and systematic training to identify and utilize relevant information.
 This gap presents a challenge in enriching domain knowledge and enhancing data augmentation.
-Knowledge augmentation is vital for improving model performance, especially in domains with limited information or complex data structures (Tang et al.
-2020).
-In knowledge augmentation, knowledge transfer is an important method that improves learning performance by transferring knowledge from external information sources.
-In data-limited scenarios, effective knowledge sourcing can bridge domain information gaps and enhance model robustness by leveraging relevant information from external sources.
-In these scenarios, knowledge augmentation, especially knowledge transfer can reduce reliance on extensive target domain data by strategically selecting sources matching the target needs.
-Despite its potential, traditional knowledge augmentation methods often involve manual intervention, where human experts pick source domains based on their subjective interpretation of the target domain’s requirements.
-Using tabular learning as an example highlights several challenges: (1) structural and format differences across domains hinder data alignment and integration; (2) discrepancies in tasks and content between the selected source and target domains can diminish the success of knowledge augmentation; (3) extensive data preprocessing is often required to match the chosen source domain dataset properly.
-Additionally, relying on human expertise for source domain selection can lead to biases and inefficiencies, as these choices are typically made based on prior knowledge rather than through a rigorous, data-driven analysis.
-A natural idea is to construct an automated search for datasets with relevant knowledge in a database with a similar structure to the target data, such as utilizing Retrieval-Augmented Generation (RAG).
-RAG combines retrieval systems with generative language models, and it enhances the model’s capabilities by providing access to an external library (Guo et al.
-2017; Song et al.
-2019; Shi et al.
-2018).
-For dataset retrieval, some existing works (Fleischer et al.
-2024; Siriwardhana et al.
-2023) use the retrieved documents as contextual information for generation and significantly enrich the model’s knowledge base.
-However, this direct method often proves impractical due to the high computational costs of embedding entire domains or whole datasets.
-Additionally, the retrieved knowledge may not align effectively with the intended application.
-Such misalignments typically result from variations in data distribution, feature spaces, or contextual differences between the source and target data.
-Furthermore, the quality of the retrieved data, such as the noise or incomplete features, can also hinder its effectiveness for augmenting data via knowledge augmentation.
-Thus, in such situations, identifying the optimal knowledge source that provides the most relevant and high-quality data tailored to the specific needs of the target data through the capabilities of LLMs remains a substantial challenge.
+Knowledge augmentation is vital for improving model performance, especially in domains with limited information or complex data structures Tang et al.
+(2020).
+In knowledge augmentation, knowledge transfer is an important method improving learning performance by transferring knowledge from external information sources Khodaee et al.
+(2024).
+In data-limited scenarios, effective knowledge sourcing can bridge domain information gaps and enhance model robustness using relevant external information.
+In these scenarios, knowledge augmentation, especially knowledge transfer, can reduce reliance on extensive target domain data by strategically selecting sources matching target needs.
+Despite its potential, traditional knowledge augmentation methods often involve manual intervention Ringwald and Stiefelhagen (2021); Nam et al.
+(2024); Zhang et al.
+(2024d), where human experts select source domains based on their subjective interpretation of the target domain’s requirements.
+Using tabular learning as an example highlights several challenges: (1) structural and format differences across domains hinder data alignment and integration; (2) discrepancies in tasks and content between the selected source and target domains can diminish the success of knowledge augmentation; and (3) extensive data preprocessing is often required to properly match the chosen source domain dataset.
+Additionally, relying on human expertise for source domain selection can lead to biases and inefficiencies, as these choices are typically made based on prior knowledge rather than a rigorous, data-driven analysis.
+A natural idea is to construct an automated search for datasets with relevant knowledge in a database that has a similar structure to the target data, such as utilizing Retrieval-Augmented Generation (RAG) Gao et al.
+(2023); Lewis et al.
+(2020).
+RAG combines retrieval systems with generative language models, and it enhances the model’s capabilities by providing access to an external library Guo et al.
+(2017); Zhang et al.
+(2025b); Shi et al.
+(2018).
+For dataset retrieval, some existing works Fleischer et al.
+(2024); Siriwardhana et al.
+(2023) use the retrieved documents as contextual information for generation and significantly enrich the model’s knowledge base.
+However, this direct method is often impractical due to the high costs of embedding entire domains or datasetsSeemakhupt et al.
+(2024); Jin et al.
+(2024).
+Additionally, retrieved knowledge may not align effectively with the intended applicationEdge et al.
+(2024); Li and Ramakrishnan (2025).
+Such misalignments typically result from variations in data distribution, feature spaces, or contextual differences between source and target data.
+Furthermore, retrieved data quality, like noise or incomplete features, can also hinder its effectiveness for augmenting data via knowledge augmentation.
+Thus, in such situations, identifying the optimal knowledge source providing the most relevant, high-quality data tailored to specific target data needs via LLM capabilities remains a substantial challenge.
 Figure 1: Example of LEKA.
@@
 Our Targets.
-Comprehensively, we aim to address three main challenges in retrieving source domain data: (1) how to effectively extract key information based on the target domain with limited computational costs; (2) how to design an efficient automatic source domain retrieval for knowledge augmentation; and (3), how to automatically harmonize the retrieved source data with the target domain to optimize transfer efficiency and improve learning outcomes.
+We aim to address three main challenges in retrieving source domain data: (1) how to effectively extract key information from the target domain with limited computational cost; (2) how to design an efficient, automatic source domain retrieval for knowledge augmentation; and (3) how to automatically harmonize the retrieved source data with the target domain to optimize transfer efficiency and improve learning outcomes.
 Our Approach.
 To address the challenges above, we design LLM-Enhanced Knowledge Augmentation (LEKA), a novel and automated data retrieval and augmentation method by knowledge augmentation.
-Specifically, (1) we utilize an LLM to extract the key textual information of the target domain; (2) we deploy dataset RAG in an external database to efficiently extract relevant knowledge to target domain dataset.
+Specifically, (1) we utilize an LLM to extract the key textual information of the target domain; (2) we deploy dataset-RAG in an external database to efficiently extract knowledge relevant to the target domain dataset.
 The RAG libraries can contain large amounts of continuously updated datasets, ensuring that the data retrieved from these libraries is more precise and timely; (3) then the LLM automatically harmonizes the retrieved source data with the target domain on feature space and marginal probability measures to enhance downstream task learning performance.
@@
 This data harmonization is crucial for reducing domain shift, which optimizes transfer efficiency and improves the overall learning outcomes of the model.
-As shown in Figure 1, for a target domain targeting predicting a rare cancer, direct learning is challenging due to insufficient sample information.
-To augment the data with external information, we need human experts with knowledge of both cancer physiological indicators and machine learning to retrieve datasets.
-However, our method can automate the retrieval of a common cancer dataset as the knowledge source with extensive feature information.
-This retrieved dataset has a more complete and transparent feature space with a structure similar to the target domain.
+As shown in Figure 1, consider a target domain focused on predicting a rare cancer, direct learning is challenging due to insufficient sample information.
+To augment the data with external information, would typically require human experts with knowledge of both cancer physiological indicators and machine learning to retrieve datasets.
+In contrast, our method can automate the retrieval of a common cancer dataset as the knowledge source with extensive feature information.
+This retrieved dataset possesses a more complete and transparent feature space with a structure similar to the target domain.
 The LLM then adjusts the retrieved data based on the characteristics of the target domain.
@@
 Experimental results demonstrate that our method has clear advantages over existing methods.
-Related Work
-Knowledge Transfer
-Knowledge transfer is a knowledge augmentation method that improves learning on new tasks by transferring knowledge from a related task (Alyafeai, AlShaibani, and Ahmad 2020; Iman, Arabnia, and Rasheed 2023; Bu et al.
-2021).
+2 Related Work
+2.1 Knowledge Transfer
+Knowledge transfer is a knowledge augmentation method that improves learning on new tasks by transferring knowledge from a related task Alyafeai et al.
+(2020); Wang et al.
+(2022, 2025).
 Knowledge transfer allows cross-domain knowledge transformation despite data distribution or feature space differences.
-Specifically, knowledge transfer adapts models developed for one task to perform better on different but related tasks, as it adjusts the feature mappings and decision boundaries to suit new knowledge (He et al.
-2021; Han, Pang, and Wu 2021; Yordanov et al.
-2021).
-Finding suitable pre-training knowledge is crucial for knowledge transfer because it significantly enhances the effectiveness of the model by providing a relevant starting point.
+Specifically, knowledge transfer adapts models developed for one task to perform better on different but related tasks, as it adjusts the feature mappings and decision boundaries to suit new knowledge He et al.
+(2021); Han et al.
+(2021); Yordanov et al.
+(2021).
+Finding suitable pre-training knowledge is crucial for knowledge transfer because it significantly enhances the model’s effectiveness by providing a relevant starting point.
 While many methods exist to adjust existing knowledge within the same or across different domains to improve transfer learning outcomes, the process of retrieving certain knowledge still heavily relies on manual effort.
-Data Harmonizing with LLMs
-Data harmonization using LLMs offers significant benefits by leveraging their natural language capabilities to standardize and integrate diverse datasets and further enhance model performance through improved data consistency (Cao, Gorodetsky, and Mitkas 2009; Liu et al.
-2021; Durante et al.
-2024; Zhang et al.
-2024d).
-However, this approach has several challenges, including the high computational costs of training LLMs and potential biases in training data, which can adversely affect the harmonization process (Feng et al.
-2021; Li, Hou, and Che 2022; Bu et al.
-2019).
+2.2 Data Harmonizing with LLMs
+Data harmonization using LLMs offers significant benefits by leveraging their natural language capabilities to standardize and integrate diverse datasets and further enhance model performance through improved data consistency Cao et al.
+(2009); Zhang et al.
+(2024a); Liu et al.
+(2021); Durante et al.
+(2024).
+However, this approach has several challenges, including the high computational costs of training LLMs and potential biases in training data, which can adversely affect the harmonization process Feng et al.
+(2021); Xie et al.
+(2024); Zhang et al.
+(2025a).
 Furthermore, the risk of overfitting remains a concern, as models may become overly specialized in training data nuances, reducing their effectiveness on new datasets.
-Retrieval Augmented Generation for Knowledge Augmentation
-Retrieval Augmented Generation (RAG) (Lewis et al.
-2020) for knowledge augmentation is a method that enhances the capabilities of generative language models by integrating information retrieval with model generation (Hu and Lu 2024).
-This technique aids LLMs in tasks that demand deep and specific domain knowledge (Zhang et al.
-2024b; Huang and Huang 2024; Zhang et al.
-2024a).
-RAG for knowledge augmentation provides access to expansive external libraries (collections of documents or domain-relevant knowledge), making it especially effective for transferring knowledge across different domains (Siriwardhana et al.
-2023; Dan, Lu, and Li 2024).
+2.3 Retrieval Augmented Generation for Knowledge Augmentation
+Retrieval Augmented Generation (RAG) Lewis et al.
+(2020) for knowledge augmentation is a method that enhances the capabilities of generative language models by integrating information retrieval with model generation Hu and Lu (2024).
+This technique aids LLMs in tasks that demand deep and specific domain knowledge Zhang et al.
+(2024b); Huang and Huang (2024).
+RAG for knowledge augmentation provides access to expansive external libraries (collections of documents or domain-relevant knowledge), making it especially effective for transferring knowledge across different domains Siriwardhana et al.
+(2023).
 RAG for knowledge augmentation embeds extensive databases directly into the generative process so that specific, domain-related information is both accessible and effectively utilized.
 In this way it can significantly enhance the performance for complex tasks in knowledge augmentation scenarios.
-Preliminary
-Definition
+3 Preliminary
+3.1 Definition
 Definition 1. (Domain)
@@
 Source Domain.
-The source domain $\mathcal{D}_{S}$ consists of instances paired with labels $y$, represented as $\mathcal{D}_{S}=\{(\mathbf{x},y)|\mathbf{x}_{i}\in\mathcal{X}_{S},y_{i}\in\mathcal{Y}_{S},i=1,\ldots,n^{S}\}$.
+The source domain $\mathcal{D}_{S}$ consists of instances paired with labels $y$, represented as $\mathcal{D}_{S}=\{(\mathbf{x},\mathbf{y})|\mathbf{x}_{i}\in\mathcal{X}_{S},y_{i}\in\mathcal{Y}_{S},i=1,\ldots,n^{S}\}$.
 Here, $\mathcal{X}_{S}$ is the feature space and $\mathcal{Y}_{S}$ is the label space for the source domain.
@@
 Target Domain.
-The target domain $\mathcal{D}_{T}$ typically contains a mix of unlabeled instances and a smaller set of labeled instances, denoted as $\mathcal{D}_{T}=\{\mathbf{x}\in\mathcal{X}_{T}\}\cup\{(\mathbf{x},y)|(\mathbf{x},y)\in\mathcal{X}_{T}\times\mathcal{Y}_{T}\}$.
+The target domain $\mathcal{D}_{T}$ typically contains a mix of unlabeled instances and a smaller set of labeled instances, denoted as $\mathcal{D}_{T}=\{\mathbf{x}\in\mathcal{X}_{T}\}\cup\{(\mathbf{x},\mathbf{y})|(\mathbf{x},\mathbf{y})\in\mathcal{X}_{T}\times\mathcal{Y}_{T}\}$.
 Here, $\mathcal{X}_{T}$ is the feature space and $\mathcal{Y}_{T}$ is the label space for the target domain.
@@
 The goal of knowledge transfer is to utilize the knowledge embedded in the source domains $\mathcal{D}_{S_{i}}$ to enhance the performance of the learned decision functions $f_{T_{j}}$ across the target domains $\mathcal{D}_{T_{j}}$, for $j=1,\ldots,m^{T}$.
-Problem Formulation
+3.2 Problem Formulation
 We formulate the task to enhance the performance of knowledge transfer through automated data retrieval and harmonization using an LLM.
@@
 where $\mathcal{P}$ is the performance indicator of $f_{T_{j}}$ and $\hat{\mathcal{D}_{S}}$ is a reconstructed source domain aligned with $\mathcal{D}_{T}$.
-Methodology
-Figure 2: Framework of LEKA includes: 1) an LLM extracts and embeds the textual information of target dataset, then 2) retrieves datasets in libraries and 3) process data harmonization.
-With harmonized dataset, we can transfer knowledge from the source dataset we construct to enhance target dataset learning.
+4 Methodology
+Figure 2: Framework of LEKA includes: 1) an LLM extracts and embeds the textual information of the target dataset, then 2) retrieves datasets in libraries and 3) processes data harmonization.
+With harmonized datasets, we can transfer knowledge from the source dataset we construct to enhance learning on the target dataset.
 In this section, we introduce our novel knowledge augmentation method named LLM-Enhanced Knowledge Augmentation (LEKA), designed to dynamically and automatically retrieve and refine source data to transfer knowledge and further enhance target data learning.
@@
 Figure 2 illustrates the overview of the LEKA framework, which comprises three stages: (1) Dataset Retrieval; (2) Data Harmonization; and (3) Knowledge Transfer and Evaluation.
-Dataset Retrieval
+4.1 Dataset Retrieval
 In this phase, our primary objective is to identify and retrieve a source dataset $\mathcal{D}_{S}$ similar to the target dataset $\mathcal{D}_{T}$ for effective pre-training in knowledge transfer scenarios.
@@
 However, this approach can be computationally intensive and prone to errors.
-Inspired by the strategies outlined in (Zhang et al.
-2024c), we turn to leverage the textual information, including the data descriptions $\mathcal{C}_{T}$ and feature names $l_{T}$ of $\mathcal{D}_{T}$.
+Inspired by the strategies outlined in Zhang et al.
+(2024c), we turn to leverage the textual information, including the data descriptions $\mathcal{C}_{T}$ and feature names $l_{T}$ of $\mathcal{D}_{T}$.
 These textual elements provide rich insights into the structure and purpose of $\mathcal{D}_{T}$, and the LLM can precisely focuses on the semantic content of $\mathcal{D}_{T}$.
@@
 where $\text{embed}(\cdot)$ is the embedding process that transforms the input into a vector representation.
-We then retrieve through in a library $\mathcal{L}$ of dataset for the top-$k$ most relevant datasets $\{\mathcal{D}_{k}\}$ to query $\mathcal{Q}_{t}$.
+We then retrieve from a library $\mathcal{L}$ of datasets for the top-$k$ most relevant datasets $\{\mathcal{D}_{k}\}$ to query $\mathcal{Q}$.
 We evaluate the relevance of all datasets in $\mathcal{L}$ to the query $\mathcal{Q}$ by the cosine similarity of the embeddings and select the top-$k$ datasets with the highest similarity:
@@
 This prepares us for further dataset processing for transfer learning.
-Data Harmonization
-Now we have $k$ potential source domain datasets, we turn to align these datasets with the target dataset $\mathcal{D}_{T}$ with an LLM.
+4.2 Data Harmonization
+Now that we have $k$ potential source domain datasets, we turn to align these datasets with the target dataset $\mathcal{D}_{T}$ with an LLM.
 This alignment process is to adjust the feature space and marginal probability measures of the source dataset $\mathcal{D}_{k}$ to closely match those of the target dataset $\mathcal{D}_{T}$.
@@
 Feature Space Transformation.
-We denote $\mathcal{X}_{k}$ and $\mathcal{X}_{T}$ to be the feature spaces of the source dataset $\mathcal{D}_{k}$ and the target dataset $\mathcal{D}_{T}$, respectively.
-Our LLM constructs a mapping function $f_{map}:\mathcal{X}_{k}\rightarrow\mathcal{X}_{T}$ for alignment.
+We denote $\mathcal{X}_{k}$ and $\mathcal{X}_{T}$ as the feature spaces of the source dataset $\mathcal{D}_{k}$ and the target dataset $\mathcal{D}_{T}$, respectively.
+Our LLM constructs a mapping function $f:\mathcal{X}_{k}\rightarrow\mathcal{X}_{T}$ for alignment.
 This function transforms the features in $\mathcal{X}_{k}$ to minimize the distance between the transformed source features and the target features.
@@
 Our LLM then aligns the marginal probability measures $\mathcal{P}_{k}$ and $\mathcal{P}_{T}$ of the source and target datasets.
-The LLM analyzes and refines the textual information including labels, feature names, and classification probability distribution,.etc.
+The LLM analyzes and refines the textual information, including labels, feature names, and classification probability distribution, etc.
 To measure the differences between $\mathcal{D}_{k}$ and $\mathcal{D}_{T}$, we adopt the Wasserstein distance:
@@
 With this source-target alignment process, the LLM improves the efficiency of knowledge transfer from the source domain to the target domain.
-Thus it enhances the overall performance of our transfer learning models.
-Further, we optimize data handling and boost the adaptability and accuracy of transfer learning to be adopted in various applications.
-Knowledge Transfer and Evaluation
+Thus, it enhances the overall performance of our transfer learning models.
+Further, we optimize data handling and boost the adaptability and accuracy of transfer learning for adoption in various applications
+4.3 Knowledge Transfer and Evaluation
 After aligning the source dataset $\mathcal{D}_{k}^{\prime}$ with the target dataset $\mathcal{D}_{T}$, and adjusting the marginal probability measures $\mathcal{P}_{k}^{\prime}$ to $\mathcal{P}_{T}$, we proceed to integrate these aligned datasets into the transfer learning model.
@@
 We calculate the expectation by the probability distribution $\mathcal{P}_{T}$, which has been closely aligned with $\mathcal{P}_{k}^{\prime}$ to ensure consistency and maximize the efficacy of the knowledge transfer.
-During update, the optimization of model parameters $\theta$ leverages both the target dataset $\mathcal{D}_{T}$ and the aligned source dataset $\mathcal{D}_{k}^{\prime}$, which incorporates domain-specific characteristics to reduce domain shift.
+During update, the optimization of model parameters $\theta$ leverages both the target dataset $\mathcal{D}_{T}$ and the aligned source dataset $\mathcal{D}_{k}^{\prime}$, a dataset that incorporates domain-specific characteristics to reduce domain shift.
 We define the optimization process as:
@@
 The expectations $\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}$ and $\mathbb{E}_{(x,y)\in\mathcal{D}_{T}}$ represent the mean losses over the source and target datasets, respectively.
-The weighting factor $\alpha$ adjusts the relative influence of each dataset, and keeps flexible adaptation between leveraging established knowledge from the source and integrating new data from the target.
+The weighting factor $\alpha$ adjusts the relative influence of each dataset and enables flexible adaptation between leveraging established knowledge from the source and integrating new data from the target.
 In this way, we utilize the transferred knowledge to enhance model performance on the target task.
@@
 In this way, we systematically improve the model based on empirical performance metrics.
-Table 1: Datasets description. Here we use four datasets from medical and economic domains.
 Datasets Samples Features Class
@@
 TCC 7043 21 2
-Experiments
+Table 1: Datasets description. Here we use four datasets from medical and economic domains.
+5 Experiments
 In this section, we present four experiments to demonstrate the effectiveness and impacts of the LEKA.
 First, we compare the performance of the LEKA against several baseline methods on four downstream tasks.
-Second, we present the correlations between several target domains and their retrieved source domain.
+Second, we present the correlations between several target domains and their retrieved source domains.
 Finally, we discuss the reason for the improvement of performance.
-Table 2: Performance comparison of transfer learning methods.
 Dataset Metrics FTT TTab LEKA
@@
 F1 0.817 0.722 0.865
+Table 2: Performance comparison of transfer learning methods.
 Figure 3: Comparison of accuracy and F1 scores on various transfer learning methods.
-Experiment Settings
+5.1 Experiment Settings
 Datasets and Domains.
-We evaluate our method primarily on four datasets of medical and economic domains: (1) Breast Cancer Wisconsin (Diagnostic) (BCW) (Wolberg et al.
-1995), (2) Heart Disease (HD) (Janosi et al.
-1988), (3) Vehicle Insurance Data (VID) (Bhatt 2019), and (4) Telco Customer Churn (TCC) (Zhang, Moro, and Ramos 2022).
+We evaluate our method primarily on four datasets of medical and economic domains: (1) Breast Cancer Wisconsin (Diagnostic) (BCW) Wolberg et al.
+(1995), (2) Heart Disease (HD) Janosi et al.
+(1989), (3) Vehicle Insurance Data (VID) Bhatt (2019), and (4) Telco Customer Churn (TCC) BlastChar (2018).
 We show the detailed information about features of the datasets in Table 1.
@@
 Recall (Rec), also known as sensitivity, reflects the ratio of true positive predictions to all actual positives for each class.
-F-Measure (F1) is the harmonic mean of precision and recall, providing a single score that balances both metrics.
-We apply the LEKA across a range of models: 1) Tabnet (TN) (Arik and Pfister 2021); 2) TabTransformer (TT) (Huang et al.
-2020); 3) Random Forest (RF) (Rigatti 2017); 4) Gradient Boosting Decision Trees (GBDT) (Lin et al.
-2023) ; 5) XGBoost (XB) (Chen and Guestrin 2016).
-We compare the performance outcomes in these tasks both with and without our method.
+F-Measure (F1) is the harmonic mean of precision and recall, calculated here as the macro-average.
+We apply the LEKA across a range of models: 1) Tabnet (TN) Arik and Pfister (2021); 2) TabTransformer (TT) Huang et al.
+(2020); 3) Random Forest (RF) Rigatti (2017); 4) Gradient Boosting Decision Trees (GBDT) Lin et al.
+(2023) ; 5) XGBoost (XB) Chen and Guestrin (2016).
+We compare the performance in these tasks both with and without our method.
 Baseline Models.
-We compare the LEKA method with four baseline methods, including: 1) Raw Data: direct using vanilla data; 2) TIFG (Zhang et al.
-2024c): feature generation with LLM; 3) KPDDS (Huang et al.
-2024): data synthesis with LLM using key points examples; 4) GReaT (Borisov et al.
-2022): data synthesis with LLM simulating data subset distributions; 5) FTT (Levin et al.
-2022) A TabTransformer pretrained on the source domain and then fine-tuned on the target domain; 6) TTab (Wang and Sun 2022) A transferable tabular Transformer capable of learning from multiple tabular datasets.
+We compare the LEKA with six baseline methods, including: 1) Raw Data: using vanilla data; 2) TIFG Zhang et al.
+(2024c): feature generation with LLM; 3) KPDDS Huang et al.
+(2024): data synthesis with LLM using key point examples; 4) GReaT Borisov et al.
+(2022): data synthesis with LLM simulating data subset distributions; 5) FTT Levin et al.
+(2022): A TabTransformer pretrained on the source domain and then fine-tuned on the target domain; 6) TTab Wang and Sun (2022): A transferable tabular Transformer capable of learning from multiple tabular datasets.
 Implementation Details.
@@
 For the RF and GBDT models, the number of trees is set to $100$, with GBDT also configured with a learning rate of $0.1$ and a max depth of $3$.
-TTab is set with a maximum epoch of $50$, a learning rate of $1\times 10^{-3}$, and a weight decay of $1\times 10^{-4}$.
-Table 3: Overall performance on downstream tasks. The best results are highlighted in bold, and the runner-up results are highlighted in underline. (Higher values indicate better performance.)
+TTab is set with a maximum of $50$ epochs, a learning rate of $1\times 10^{-3}$, and a weight decay of $1\times 10^{-4}$.
 Metrics Model GBDT RF TN XB TT
@@
 LEKA 0.990 0.996 0.866 0.756 0.982 0.864 0.885 0.739 0.981 0.568 0.916 0.760 0.991 0.959 0.885 0.761 0.991 0.780 0.750 0.865
-Experiment Results
-Overall Performance. The experiment results are shown in Table 3, where we compare our LEKA method in four datasets. In summary:
+Table 3: Overall performance on downstream tasks. The best results are highlighted in bold, and the runner-up results are highlighted in underline. (Higher values indicate better performance.)
+5.2 Experiment Results
+Overall Performance.
+The experiment results are shown in Table 3, where we compare our LEKA method across four datasets. In summary:
 (1) Compared with baseline models, LEKA significantly improves accuracy, surpassing baseline methods like TIFG and GReaT, with enhancements up to $4.4\%$ in models like GBDT.
@@
 The F1 scores under LEKA are notably high, reflecting its effective balance between precision and recall across various models and datasets.
-These results validate the effectiveness of LEKA ’s retrieval and harmonization strategies and the robustness in diverse application scenarios.
+These results validate the effectiveness of LEKA ’s retrieval and harmonization strategies and its robustness in diverse application scenarios.
 Overall, LEKA’s strategic approach to knowledge transfer is particularly advantageous in complex data environments, showcasing its adaptability and efficiency compared to traditional methods.
@@
 For improvement of overall accuracy, we demonstrate the LEKA’s effective retrieval and harmonization of feature space and marginal probability measures with the target datasets.
+LEKA demonstrates exceptional precision, improving the TT metric for the TCC dataset by over $10\%$ compared to baselines, effectively reducing false positives.
 Meanwhile, the LEKA consistently outperforms baselines in reducing misclassification rates with the highest F1 scores.
@@
 This superior performance is consistent across other datasets like VID, HD, and TCC, with notable enhancements in precision and recall, emphasizing LEKA’s effective data harmonization capabilities.
-This adaptability and efficiency in handling diverse datasets underscore LEKA’s robustness and its significant advantages.
-Conclusion
+This adaptability and efficiency in handling diverse datasets underscore LEKA’s robustness and advantages.
+6 Conclusion
 In this paper, we introduce LLM-Enhanced Knowledge Augmentation (LEKA), a novel retrieval and harmonization framework that dynamically refines source data for effective knowledge transfer.
@@
 For future work, we plan to extend this framework to include more varied data types and conduct a thorough empirical analysis to gain further insights into its underlying mechanisms and impacts.
-Limitations
-We acknowledge the following limitations of our work: (1) The current work is only tested on tabular data.
-More complex test scenarios are not involved; (2) Despite its automation, LEKA can be computationally intensive.
+7 Limitations
+We acknowledge the following limitations of our work: (1) The current work has only been tested on tabular data, and more complex test scenarios have not yet been involved; (2) Despite its automation, LEKA can be computationally intensive.
 This could limit its applicability in resource-constrained environments or require substantial computational resources to maintain operational efficiency; (3) Adopting the LEKA in tasks with unique requirements may have inherent limitations.
+Acknowledgements
+Dr.
+Kunpeng Liu is supported by the National Science Foundation (NSF) via the grant numbers 2426339 and 2348485.
+Dr.
+Yanjie Fu is supported by the National Science Foundation (NSF) via the grant numbers 2426340, 2416727, 2421864, 2421865, 2421803, and National Academy of Engineering Grainger Foundation Frontiers of Engineering Grants.

### v3

@@
 Humans excel in analogical learning and knowledge transfer and, more importantly, possess a unique understanding of identifying appropriate sources of knowledge.
-From a model’s perspective, this presents an interesting challenge.
-If models could autonomously retrieve knowledge useful for transfer or decision-making to solve problems, they would transition from passively acquiring to actively accessing and learning from knowledge.
+From a model’s perspective, this presents a unique challenge.
+If models could autonomously retrieve knowledge relevant for transfer or decision-making to solve problems, they would transition from passively acquiring to actively accessing and learning from knowledge.
 However, filling models with knowledge is relatively straightforward—it simply requires more training and accessible knowledge bases.
@@
 This LEKA method extracts key information from the target domain’s textual information, retrieves pertinent data from external data libraries, and harmonizes retrieved data with the target domain data in feature space and marginal probability measures.
-We validate the effectiveness of our approach through extensive experiments across various domains and demonstrate significant improvements over traditional methods in reducing computational costs, automating data alignment, and optimizing transfer learning outcomes.
+We validate the effectiveness of our approach through extensive experiments across various domains and demonstrate significant improvements over traditional methods in automating data alignment and optimizing transfer learning outcomes.
 1 Introduction
-Humans are good at identifying proper sources of knowledge.
+Humans are good at identifying relevant sources of knowledge.
 This is an ability that is rooted in our capability for analogical reasoning and knowledge management.
@@
 (2020).
-In knowledge augmentation, knowledge transfer is an important method improving learning performance by transferring knowledge from external information sources Khodaee et al.
+Knowledge transfer is a crucial method within knowledge augmentation for improving learning performance by transferring knowledge from external information sources Khodaee et al.
 (2024).
@@
 We adopt an LLM to retrieve proper source domain data to transfer knowledge to a data-limited target domain.
-The LLM extracts the key information of target data to retrieve for a relevant dataset; then we adopt the LLM for harmonization.
+The LLM extracts the key information of the target data to retrieve a relevant dataset; then, we adopt the LLM for harmonization.
 Our Targets.
@@
 Specifically, (1) we utilize an LLM to extract the key textual information of the target domain; (2) we deploy dataset-RAG in an external database to efficiently extract knowledge relevant to the target domain dataset.
-The RAG libraries can contain large amounts of continuously updated datasets, ensuring that the data retrieved from these libraries is more precise and timely; (3) then the LLM automatically harmonizes the retrieved source data with the target domain on feature space and marginal probability measures to enhance downstream task learning performance.
+The RAG libraries can contain large amounts of continuously updated datasets, ensuring that the data retrieved from these libraries is more precise and timely; (3) then the LLM automatically harmonizes the retrieved source data with the target domain in feature space and marginal probability measures to enhance downstream task learning performance.
 The data harmonization reduces the structural and semantic differences between the source and target data.
 This data harmonization is crucial for reducing domain shift, which optimizes transfer efficiency and improves the overall learning outcomes of the model.
-As shown in Figure 1, consider a target domain focused on predicting a rare cancer, direct learning is challenging due to insufficient sample information.
-To augment the data with external information, would typically require human experts with knowledge of both cancer physiological indicators and machine learning to retrieve datasets.
+As shown in Figure 1, consider a target domain focused on predicting a rare cancer, so direct learning is challenging due to insufficient sample information.
+To augment the data with external information would typically require human experts with knowledge of both cancer physiological indicators and machine learning to retrieve datasets.
 In contrast, our method can automate the retrieval of a common cancer dataset as the knowledge source with extensive feature information.
@@
 Knowledge transfer allows cross-domain knowledge transformation despite data distribution or feature space differences.
-Specifically, knowledge transfer adapts models developed for one task to perform better on different but related tasks, as it adjusts the feature mappings and decision boundaries to suit new knowledge He et al.
-(2021); Han et al.
+Specifically, knowledge transfer adapts models developed for one task to perform better on different but related tasks, as it adjusts the feature mappings and decision boundaries to suit new knowledge Han et al.
 (2021); Yordanov et al.
@@
 2.2 Data Harmonizing with LLMs
-Data harmonization using LLMs offers significant benefits by leveraging their natural language capabilities to standardize and integrate diverse datasets and further enhance model performance through improved data consistency Cao et al.
+Data harmonization applying LLMs offers significant benefits by leveraging their natural language capabilities to standardize and integrate diverse datasets and further enhance model performance through improved data consistency Cao et al.
 (2009); Zhang et al.
@@
 RAG for knowledge augmentation embeds extensive databases directly into the generative process so that specific, domain-related information is both accessible and effectively utilized.
-In this way it can significantly enhance the performance for complex tasks in knowledge augmentation scenarios.
+In this way, it can significantly enhance the performance for complex tasks in knowledge augmentation scenarios.
 3 Preliminary
@@
 4 Methodology
-Figure 2: Framework of LEKA includes: 1) an LLM extracts and embeds the textual information of the target dataset, then 2) retrieves datasets in libraries and 3) processes data harmonization.
+Figure 2: Framework of LEKA includes: 1) an LLM extracts and embeds the textual information of the target dataset, then 2) retrieves datasets in libraries, and 3) processes data harmonization.
 With harmonized datasets, we can transfer knowledge from the source dataset we construct to enhance learning on the target dataset.
@@
 (2024c), we turn to leverage the textual information, including the data descriptions $\mathcal{C}_{T}$ and feature names $l_{T}$ of $\mathcal{D}_{T}$.
-These textual elements provide rich insights into the structure and purpose of $\mathcal{D}_{T}$, and the LLM can precisely focuses on the semantic content of $\mathcal{D}_{T}$.
+These textual elements provide insights into the structure and purpose of $\mathcal{D}_{T}$, and the LLM can precisely focus on the semantic content of $\mathcal{D}_{T}$.
 First, the LLM $\mathcal{A}$ constructs a query $\mathcal{Q}$ based on this textual information:
@@
 Now that we have $k$ potential source domain datasets, we turn to align these datasets with the target dataset $\mathcal{D}_{T}$ with an LLM.
-This alignment process is to adjust the feature space and marginal probability measures of the source dataset $\mathcal{D}_{k}$ to closely match those of the target dataset $\mathcal{D}_{T}$.
+This alignment process adjusts the feature space and marginal probability measures of the source dataset $\mathcal{D}_{k}$ to closely match those of the target dataset $\mathcal{D}_{T}$.
 This process involves transformations of features and adjustments of distributions.
@@
 We adopt a Gaussian kernel $\kappa(x,y)=\exp(-\gamma\|x-y\|^{2})$, with $\gamma$ being a positive bandwidth parameter.
-This kernel function can effectively measure the similarity between pairs of points in high-dimensional spaces and captures both linear and non-linear relationships.
-Thus, it is capable of handling the complexities inherent in high-dimensional data.
-Specifically, compared with Euclidean distances, kernel distances can capture the geometric structure of the data manifold.
-It provides a more robust and informative measure of similarity for the source-target alignment.
+This kernel function can effectively measure the similarity between points in high-dimensional spaces and captures both linear and non-linear relationships.
+Thus, it can handle the complexities inherent in high-dimensional data.
+Compared with Euclidean distances, kernel distances can capture the geometric structure of the data manifold.
+It provides a more robust and informative similarity measure for the source-target alignment.
 Marginal Probability Measures Harmonization.
-Our LLM then aligns the marginal probability measures $\mathcal{P}_{k}$ and $\mathcal{P}_{T}$ of the source and target datasets.
-The LLM analyzes and refines the textual information, including labels, feature names, and classification probability distribution, etc.
+Our LLM aligns the marginal probability measures $\mathcal{P}_{k}$ and $\mathcal{P}_{T}$ of the source and target datasets.
+The LLM analyzes and refines textual information such as labels, feature names, and classification probability distributions.
 To measure the differences between $\mathcal{D}_{k}$ and $\mathcal{D}_{T}$, we adopt the Wasserstein distance:
-[…]
+$W(\mathcal{P}_{k},\mathcal{P}_{T})=\inf_{\gamma\in\Gamma(\mathcal{P}_{k},\mathcal{P}_{T})}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|\,d\gamma(x,y),$ (6)
 $\mathcal{D}_{k}^{{}^{\prime}}=\mathcal{A}(\mathcal{D}_{k},\mathcal{P}_{k},\mathcal{P}_{T}),$ (7)
@@
 Specifically, the Wasserstein distance calculates the minimum “geographical” cost required to move data from one distribution to another, where “geographical” cost refers to the cost of moving data from one point to another in the feature space.
-Moreover, the Wasserstein distance is particularly useful when dealing with distributions whose support sets (i.e., the effective range or set of the distributions) do not fully overlap.
+Moreover, the Wasserstein distance is advantageous when dealing with distributions whose support sets (i.e., the effective range or set of the distributions) do not fully overlap.
 In real-world data applications, it is common for the source and target datasets to originate from different distributions.
 Their data points do not align perfectly and cover the same areas.
-In these cases, traditional distance metrics like the Euclidean distance may not effectively reflect the true differences between the two datasets, as they merely measure differences in position but ignore the overall structure of the data distributions.
+In these cases, traditional distance metrics like the Euclidean distance poorly reflect the actual differences between the two datasets, as they merely measure differences in position but ignore the overall structure of the data distributions.
 With this source-target alignment process, the LLM improves the efficiency of knowledge transfer from the source domain to the target domain.
 Thus, it enhances the overall performance of our transfer learning models.
-Further, we optimize data handling and boost the adaptability and accuracy of transfer learning for adoption in various applications
+Further, we optimize data handling and boost the adaptability and accuracy of transfer learning for adoption in various applications.
 4.3 Knowledge Transfer and Evaluation
 After aligning the source dataset $\mathcal{D}_{k}^{\prime}$ with the target dataset $\mathcal{D}_{T}$, and adjusting the marginal probability measures $\mathcal{P}_{k}^{\prime}$ to $\mathcal{P}_{T}$, we proceed to integrate these aligned datasets into the transfer learning model.
-At this stage, our goal is to effectively transfer knowledge from $\mathcal{D}_{k}^{\prime}$ to $\mathcal{D}_{T}$ by minimizing a defined loss function while adapting the model to the target domain.
+At this stage, our goal is to transfer knowledge from $\mathcal{D}_{k}^{\prime}$ to $\mathcal{D}_{T}$ by minimizing a defined loss function while adapting the model to the target domain.
 To achieve this, the transfer learning process focuses on updating the decision function $f_{T_{j}}$.
@@
 We calculate the expectation by the probability distribution $\mathcal{P}_{T}$, which has been closely aligned with $\mathcal{P}_{k}^{\prime}$ to ensure consistency and maximize the efficacy of the knowledge transfer.
-During update, the optimization of model parameters $\theta$ leverages both the target dataset $\mathcal{D}_{T}$ and the aligned source dataset $\mathcal{D}_{k}^{\prime}$, a dataset that incorporates domain-specific characteristics to reduce domain shift.
+During update, the optimization of model parameters $\theta$ leverages both the target dataset $\mathcal{D}_{T}$ and the aligned source dataset $\mathcal{D}_{k}^{\prime}$, incorporating domain-specific characteristics to reduce domain shift.
 We define the optimization process as:
-$\theta^{*}$ $=\arg\min_{\theta}\bigg{(}\alpha\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}[\mathcal{L}(f_{T_{j}}(x;\theta),y)]$ (9)
+$\theta^{*}$ $=\arg\min_{\theta}\bigg(\alpha\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}[\mathcal{L}(f_{T_{j}}(x;\theta),y)]$
 $\quad+(1-\alpha)\mathbb{E}_{(x,y)\in\mathcal{D}_{T}}[\mathcal{L}(f_{T_{j}}(x;\theta),y)]\bigg).$ (9)
@@
 The weighting factor $\alpha$ adjusts the relative influence of each dataset and enables flexible adaptation between leveraging established knowledge from the source and integrating new data from the target.
-In this way, we utilize the transferred knowledge to enhance model performance on the target task.
+In this way, we apply the transferred knowledge to enhance model performance on the target task.
 Following the optimization of the model parameters, the backpropagation process then updates $\theta$ by minimizing the total loss.
@@
 $\theta\leftarrow\theta-\eta\nabla_{\theta}\left(\alpha\mathcal{L}(\mathcal{D}_{k}^{\prime};\theta)+(1-\alpha)\mathcal{L}(\mathcal{D}_{T};\theta)\right),$ (10)
-where $\eta$ is the learning rate. The gradients are computed based on both datasets, which allows the model to learn both from the aligned source data and the target data.
-After the parameter optimization and backpropagation processes, we evaluate the effectiveness of the transfer learning process; we adopt a performance metric $\phi$ on the target dataset:
+where $\eta$ is the learning rate. The gradients are computed based on both datasets, which allows the model to learn from both the aligned source data and the target data.
+After the parameter optimization and backpropagation, we evaluate the effectiveness of the transfer learning process; we adopt a performance metric $\phi$ on the target dataset:
 $\phi(f_{T_{j}}^{*},\mathcal{D}_{T}),$ (11)
@@
 TCC 7043 21 2
-Table 1: Datasets description. Here we use four datasets from medical and economic domains.
+Table 1: Datasets description. Here we use four datasets from the medical and economic domains.
 5 Experiments
 In this section, we present four experiments to demonstrate the effectiveness and impacts of the LEKA.
-First, we compare the performance of the LEKA against several baseline methods on four downstream tasks.
+First, we compare the LEKA against several baseline methods on four downstream tasks.
 Second, we present the correlations between several target domains and their retrieved source domains.
-Finally, we discuss the reason for the improvement of performance.
+Finally, we discuss the reason for performance improvement.
 Dataset Metrics FTT TTab LEKA
@@
 Datasets and Domains.
-We evaluate our method primarily on four datasets of medical and economic domains: (1) Breast Cancer Wisconsin (Diagnostic) (BCW) Wolberg et al.
+We evaluate our method on four datasets of medical and economic domains: (1) Breast Cancer Wisconsin (Diagnostic) (BCW) Wolberg et al.
 (1995), (2) Heart Disease (HD) Janosi et al.
 (1989), (3) Vehicle Insurance Data (VID) Bhatt (2019), and (4) Telco Customer Churn (TCC) BlastChar (2018).
-We show the detailed information about features of the datasets in Table 1.
+We show the detailed information about the features of the datasets in Table 1.
 Metrics and Models.
@@
 Implementation Details.
-In our setup for data synthesis and model training, we utilize GPT-4o as the query generator, combined with the Exa API to fetch web pages containing datasets from Kaggle and the UCI Machine Learning Repository that may be suitable for knowledge transfer.
+In our setup for data synthesis and model training, we utilize GPT-4o OpenAI (2024) as the query generator, combined with the Exa API Exa (2024) to fetch web pages containing datasets from Kaggle Kaggle (2024) and the UCI Machine Learning Repository University of California, Irvine (2024) that may be suitable for knowledge transfer.
 We extract dataset descriptions from web pages, and GPT-4o assesses their potential for knowledge transfer.
@@
 TTab is set with a maximum of $50$ epochs, a learning rate of $1\times 10^{-3}$, and a weight decay of $1\times 10^{-4}$.
-Metrics Model GBDT RF TN XB TT
+Metrics Model GBDT RF TN XB $T$
 Acc BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC BCW VID HD TCC
@@
 Overall Performance.
-The experiment results are shown in Table 3, where we compare our LEKA method across four datasets. In summary:
+The results are shown in Table 3, comparing our LEKA across four datasets. In summary:
 (1) Compared with baseline models, LEKA significantly improves accuracy, surpassing baseline methods like TIFG and GReaT, with enhancements up to $4.4\%$ in models like GBDT.
@@
 Meanwhile, the LEKA consistently outperforms baselines in reducing misclassification rates with the highest F1 scores.
-These results prove that LEKA’s retrieval and data harmonization reduce the likelihood of misclassification by forming a deeper and clearer understanding of the potential relationship between features.
+These results prove that LEKA’s retrieval and data harmonization reduce misclassification by forming a deeper and clearer understanding of the potential relationship between features.
 Comparison with transfer learning methods.
@@
 In these scenarios, the transfer learning methods show their reliance on manual data selection and alignment processes.
-The results demonstrate that LEKA outperforms transfer learning methods FFT and TTab across various metrics and datasets.
+The results demonstrate that LEKA outperforms transfer learning methods, FFT, and TTab across various metrics and datasets.
 For example, in the BCW dataset, LEKA improves accuracy by $3.5\%$ and recall by $3.3\%$ compared to the competing methods.
@@
 In this paper, we introduce LLM-Enhanced Knowledge Augmentation (LEKA), a novel retrieval and harmonization framework that dynamically refines source data for effective knowledge transfer.
-This structure significantly enhances the process of data augmentation by leveraging advanced LLM capabilities to automatically align and optimize data retrieved from diverse external libraries.
-Our extensive experiments across various tasks demonstrate the effectiveness and superiority over existing methods, especially in improving model adaptability and accuracy in complex data environments.
+This structure significantly enhances data augmentation by leveraging advanced LLM capabilities to automatically align and optimize data retrieved from diverse external libraries.
+Extensive experiments across various tasks demonstrate superiority to existing methods, especially in improving model adaptability and accuracy in complex data environments.
 By adopting LEKA on data-scarce domains, we achieve substantial improvements in learning performance and domain-specific task accuracy.
-For future work, we plan to extend this framework to include more varied data types and conduct a thorough empirical analysis to gain further insights into its underlying mechanisms and impacts.
+For future work, we plan to extend this framework to include more varied data types and conduct a thorough empirical analysis to understand its underlying mechanisms and impacts.
 7 Limitations
-We acknowledge the following limitations of our work: (1) The current work has only been tested on tabular data, and more complex test scenarios have not yet been involved; (2) Despite its automation, LEKA can be computationally intensive.
+We acknowledge the following limitations: (1) the current work has only been tested on tabular data, and more complex test scenarios have not yet been involved; (2) despite its automation, LEKA can be computationally intensive.
 This could limit its applicability in resource-constrained environments or require substantial computational resources to maintain operational efficiency; (3) adopting the LEKA in tasks with unique requirements may have inherent limitations.